### v1

# X-Reflect: Cross-Reflection Prompting for Multimodal Recommendation

Abstract
Large Language Models (LLMs) and Large Multimodal Models (LMMs) have been shown to enhance the effectiveness of enriching item descriptions, thereby improving the accuracy of recommendation systems. However, most existing approaches either rely on text-only prompting or employ basic multimodal strategies that do not fully exploit the complementary information available from both textual and visual modalities. This paper introduces a novel framework, Cross-Reflection Prompting, termed X-Reflect, designed to address these limitations by prompting LMMs to explicitly identify and reconcile supportive and conflicting information between text and images. By capturing nuanced insights from both modalities, this approach generates more comprehensive and contextually richer item representations. Extensive experiments conducted on two widely used benchmarks demonstrate that our method outperforms existing prompting baselines in downstream recommendation accuracy. Additionally, we evaluate the generalizability of our framework across different LMM backbones and the robustness of the prompting strategies, offering insights for optimization. This work underscores the importance of integrating multimodal information and presents a novel solution for improving item understanding in multimodal recommendation systems.
Introduction
Leveraging Large Language Models (LLMs) or Large Multimodal Models (LMMs) to enrich item descriptions has been shown to enhance the effectiveness of recommendation systems (Lyu et al. 2024; Xi et al. 2023; Li et al. 2023). By generating detailed and contextually relevant descriptions, LLMs and LMMs contribute to more accurate and personalized recommendations. However, many existing methods either rely on text-only prompting strategies or use basic multimodal prompting techniques that do not fully utilize the information available from both textual and visual modalities (Liu et al. 2024c; Baltrušaitis, Ahuja, and Morency 2018).
For instance, a movie described textually as a “light-hearted romantic comedy” may have a poster that, with its dark tones and serious expressions, suggests underlying tension. A text-only approach might overlook these visual cues, while a simple multimodal strategy might superficially combine the text and image without resolving their differences, leading to an incomplete or misleading representation of the item. Similarly, a grocery item labeled as “organic and healthy” might have packaging that subtly suggests indulgence, like rich colors or gourmet styling. While neither the text nor the image alone may convey the full character of the product, analyzing them together can highlight these additional dimensions, such as the balance between health and indulgence. These examples illustrate the limitations of current approaches and the need for more sophisticated methods that can effectively integrate and interpret information across modalities.
To recommend items more effectively, a deeper understanding of these items is necessary, which cannot be achieved by neglecting the rich information provided by additional modalities, such as images. Visual content often captures aspects that text alone cannot fully convey, including aesthetic qualities, emotional tone, or specific details that may significantly influence user preferences (Huang et al. 2019). Therefore, it is critical to develop methods that can seamlessly integrate and reconcile information from both text and images.
To address this need, we introduce a novel Cross-Reflection Prompting framework termed X-Reflect. This method, as illustrated in Figure 1, prompts LMMs to process textual and visual information concurrently, explicitly identifying and reconciling any supportive or conflicting elements between these modalities. By capturing the nuanced information present in both text and images, cross-reflection prompting enables the generation of more diverse and contextually rich item representations.
The contributions of this paper are as follows:
We propose a novel prompting strategy that effectively integrates and reconciles information from both textual and visual modalities, addressing the limitations of existing text-only and basic multimodal prompting approaches.
We conduct extensive experiments on two widely used recommendation benchmarks, demonstrating that our framework consistently outperforms baseline prompting methods in recommendation accuracy.
We assess the generalizability of our framework by applying it to multiple LMM backbones, showing that it remains effective across different model architectures.
In summary, this work underscores the importance of fully exploiting multimodal information in recommendation systems and presents a novel framework that enhances item understanding through a more integrated analysis of textual and visual data.
Related Work
Multimodal Recommendation System
Traditional unimodal recommendation systems, which depend on a single data source such as textual descriptions, often encounter challenges like data sparsity, cold start problems, noise, and redundancy. To mitigate these issues, researchers have developed multimodal recommendation systems that integrate diverse types of data, including text, images, audio, and video. These systems enable a more comprehensive understanding of content, as different modalities can offer varied perspectives and insights. Moreover, when one modality suffers from poor quality or sparse data, other modalities can compensate, enhancing the system’s overall robustness and reliability (Wei et al. 2019; Lyu and Luo 2022; Wang, Chen, and Zhu 2021; Huang et al. 2019; Tao et al. 2020).
Our method aligns with the goals of multimodal recommendation systems by also leveraging multiple data modalities, specifically text and images, to enrich item descriptions. However, unlike traditional multimodal approaches that may focus on designing advanced neural architectures for integrating data from different modalities, our framework introduces a novel cross-reflection prompting strategy to augment the input data.
Prompting Strategies for Recommendation
In recent years, text-only prompting strategies have gained attention for their effectiveness in recommendation systems. For example, Xi et al. (2023) introduced a method that leverages the reasoning and factual knowledge of LLMs to enhance item recommendations. Lyu et al. (2024) proposed four distinct prompting strategies designed to augment item descriptions using LLMs. While these methods have shown promise, they primarily focus on textual data and do not fully explore the potential of integrating visual information.
There has been limited exploration of multimodal prompting strategies in the context of recommendation. One notable exception is Rec-GPT4V (Liu et al. 2024c), which introduced a reasoning scheme called Visual-Summary Thought (VST). This method leverages large vision-language models (LVLMs) for multimodal recommendation by using user history as in-context user preferences and generating item image summaries to enhance the recommendation process.
Our work builds upon these existing prompting strategies. While Rec-GPT4V focuses on summarizing visual information to aid in recommendations, our cross-reflection prompting framework goes a step further by explicitly examining the relationship between textual and visual data. By doing so, our method captures nuanced information between modalities, leading to more informed and contextually rich item representations that significantly improve recommendation performance.
Cross-Reflection Prompting
In many cases, the text and image associated with a product may be consistent with each other and do not exhibit obvious contradictions, but subtle differences and complementary details can emerge when they are examined together.
This paper aims to design a framework that can more effectively leverage Large Multimodal Models to enrich item descriptions by fully exploiting information from both textual and visual modalities. To achieve this, we propose a novel approach called Cross-Reflection Prompting, referred to as X-Reflect.
The core idea of cross-reflection prompting is to provide LMMs with both textual and visual information as context, while explicitly instructing the models to identify and reconcile whether the textual and visual information supports or contradicts each other. The responses generated by the LMMs are then converted into embeddings, which serve as input item embeddings for the subsequent recommendation module, as illustrated in Figure 2.
We present two distinct strategies for implementing cross-reflection prompting: separate prompting and combined prompting.
Separate Prompting
In separate prompting, we first prompt two models, and to independently augment the text input T and image input I, respectively:
where and represent the intermediate generated responses. Next, another model, , is employed to generate a more comprehensive response based on these intermediate responses. In the prompt for , cross-reflection is incorporated as follows:
“This is the item description: {}. This is the description of the item image: {}. Evaluate whether the information presented in the item description and the image description are supportive of each other or if there are any discrepancies. If any conflicts are identified, please address these conflicts.”
Combined Prompting
In combined prompting, we use a single model, , to generate the response based on the text and image associated with the item through cross-reflection prompting:
“Please provide a summary of the item characteristics based on the provided text description and image. The item description is as follows: {text description}. Evaluate whether the information presented in the text and the image are supportive of each other or if there are any discrepancies. If any conflicts are identified, please address these conflicts.”
In both prompts, the parts related to cross-reflection are highlighted in purple. Inspired by the findings of Lyu et al. (2024), we concatenate the embeddings of these responses to construct item representations, which will be used in the subsequent recommendation module.
Remark. Similar to prior augmentation methods, cross-reflection prompting can be considered a plug-and-play module in the recommendation system, serving solely to augment item-side input, thus providing significant flexibility for the recommendation module. Moreover, although Figure 2 indicates all models as LMMs, only and need to be LMMs. and can be Large Language Models that do not take images as input. can be omitted, allowing the use of the original item description directly. This further enhances the flexibility of the entire framework. Lastly, both proprietary and open-source models can be used. We leave the investigation of incorporating supervised fine-tuning (SFT) and continuous pre-training of the models to improve the quality of responses for future work.
Experiment Setup
Datasets
Dataset | # Interactions | # Items | # Users |
MovieLens-1M | 1,000,209 | 3,706 | 6,040 |
Amazon-Software | 6,251 | 1,729 | 1,004 |
We use two widely recognized recommendation benchmarks: MovieLens-1M (Harper and Konstan 2015) and Amazon-Software (Ni, Li, and McAuley 2019). The dataset statistics are provided in Table 1.
MovieLens-1M: This dataset contains anonymous user ratings collected from MovieLens. The movie poster URLs are obtained from the movielens-poster repository.111https://github.com/babu-thomas/movielens-posters
Amazon-Software: The original dataset includes product reviews and metadata from Amazon between May 1996 and October 2018. The product metadata comprises details such as color, size, and package type, along with product images captured after users received the products. Additional metadata includes bullet-point descriptions under the product title, technical details, and tables of similar products. For our experiments, we focus on the Software category within the Amazon dataset.
Following the approach of He et al. (2017), we convert the rating data into implicit feedback, indicating whether a user has interacted with a corresponding item. In the Amazon-Software dataset, the original item descriptions are provided. However, the original MovieLens-1M data only include movie titles and genres. Therefore, we use the item descriptions for movies generated by GPT-3 from Lyu et al. (2024). For both datasets, we retain users who have rated at least five items and items that have been rated by at least five unique users.
Baselines
Our objective is to design a multimodal prompting framework that effectively enriches item descriptions by leveraging both textual and visual information for recommendation. To evaluate our cross-reflection framework, X-Reflect, we mainly compare it against two categories of baselines.
Text-only prompting baselines:
LLM-Rec (Lyu et al. 2024): This work proposes four distinct prompting strategies to enrich item descriptions. We use the recommendation-driven strategy as our baseline. The exact prompt is: “The description of an item is as follows: {item description}, what else should I say if I want to recommend it to others?”
Multimodal prompting baselines:
Rec-GPT4V (Liu et al. 2024c): This baseline prompts LMMs to generate item image summaries and integrates image comprehension in the natural language space with item titles to query user preferences. For fair comparison, we only use the instruction to generate item image summaries. The exact prompt is “What’s in this image?”
CoT (Wei et al. 2022): The Chain-of-Thought (CoT) prompting guides LLMs to generate step-by-step reasoning before delivering a final answer. Many researchers have explored its adaptation into the multimodal context (Zhang et al. 2024; Zheng et al. 2023). In our implementation, the prompt is “The item description is as follows: {item description}. Please think step by step and describe this item based on both the description and the input image. Your response should highlight how combining both the text and image provides a comprehensive understanding of the item.”
Visual-only prompting: This method prompts an LMM to describe the image: “Describe the image.”
Visual-textual prompting: It prompts an LMM to describe the item based on both the image and the original item description: “The item description is as follows: {item description}. Please describe this item based on both the description and the input image. Your response should highlight how combining both the text and image provides a comprehensive understanding of the item.”
For both categories of baselines, we convert the generated responses into embeddings using the same text encoder.
We also include three additional baselines for reference:
Item Popularity: This baseline recommends the most popular items.
Text: This baseline uses the embeddings of the original item description as input to the recommendation module.
Text + Image: This baseline uses the embeddings of the original item description combined with CLIP embeddings (Radford et al. 2021) of the associated image as input to the recommendation module.
Evaluation Protocols
We follow the evaluation methodology outlined by Wei et al. (2019). For MovieLens-1M, we split user interactions into training, validation, and test sets using an 8:1:1 ratio, while for Amazon-Software, we use a 6:2:2 ratio to ensure at least one interaction per user in the validation and test sets. Negative training samples are created by pairing users with items without recorded interactions (i.e., pseudo-negative samples). For validation and test sets, we pair each observed user-item interaction with 1,000 items the user has not interacted with. There is no overlap between negative samples in the training set and unobserved user-item pairs in the validation and test sets, ensuring the independence of evaluation data. We evaluate top-K recommendations using metrics such as Precision@K, Recall@K, and NDCG@K, where is set to 10. The reported results are averaged across five different test set splits.
Large Multimodal Models
For fair comparison, we use the same language model across all prompting strategies. For X-Reflect and other multimodal prompting baselines, we use GPT-4V (OpenAI 2023) (gpt-4-vision-preview variant). To test the effectiveness and generalizability of our framework, we also evaluate it using two open-source LMMs: LLaVA-1.5 (Liu et al. 2024a) (llava-1.5-7b-hf), an enhanced version of LLaVa (Liu et al. 2024b), and InternLM-XComposer2 (Dong et al. 2024) (internlm-xcomposer2-vl-7b), a vision-language large model based on InternLM2 (Cai et al. 2024). For text-only prompting baselines, we use GPT-4 (gpt-4 variant) (OpenAI 2023).
Implementation Details
Recommendation Module
Following Lyu et al. (2024), we evaluate different prompting baselines using a two-tower model (Wang, Zhu, and He 2021), designed to predict the likelihood of a user interacting with an item. The model consists of two towers that process user and item features into dense embeddings. The interaction likelihood is computed using the dot product of these embeddings. We use Sentence-BERT (Reimers and Gurevych 2019) to generate textual embeddings from responses, specifically using the all-MiniLM-L6-v2 model variant. These item embeddings serve as input to the item tower. The text encoder remains frozen to minimize computational costs. For the user tower, we employ an embedding table to convert user IDs into latent representations, with the output dimension set to 128 for both datasets.
Model Training
Model training is guided by binary cross-entropy loss:
where represents user-item pairs, is the set of all samples, is the label indicating interaction (1 if user has interacted with item , 0 otherwise), and is the predicted interaction likelihood. Positive samples are based on observed user-item interactions, while negative samples are generated by randomly pairing users with items they have not interacted with.
Grid Search for Hyperparameters
Model parameters are initialized randomly following a Gaussian distribution. We optimize the framework using the AdamW algorithm (Loshchilov and Hutter 2017) with a weight decay of 0.0005. Hyperparameter grids for learning rate and dropout are set to and , respectively. Early stopping is implemented with a window size and evaluation frequency of 5. The batch size is set to 4096 across all baselines. Hyperparameters that yield the highest Recall@K on the validation set are used for final testing.
Due to computational constraints, we do not perform grid search for LMM or LLM hyperparameters, instead following the settings from Lyu et al. (2024).
Computing Resources
Text generation is executed on eight NVIDIA GeForce RTX 2080 Ti GPUs, each with 11 GB of memory. Each recommendation experiment is run on a single NVIDIA GeForce RTX 1080 Ti GPU with 11 GB of memory. The open-source models are implemented using the transformers library from Hugging Face.
Results
Comparison with Baseline Prompting Strategies
Table 2 presents the average recommendation performance of X-Reflect compared to various baseline approaches. Several key observations can be made:
First, prompting methods that integrate both text and image generally outperform those relying solely on text. However, on the Amazon-Software dataset, the text-only prompting strategy LLM-Rec (Lyu et al. 2024) shows notably strong performance. The specific LLM-Rec variant implemented in this work involves instructing the LLM to describe the item with a focus on recommendation purposes. Upon closer examination, it appears that the original Amazon-Software descriptions, which contain highly technical details, may not generalize well. The responses generated through LLM-Rec, which tend to be less technical, likely contribute to its superior performance.
Second, among the prompting methods that combine text and image, those incorporating cross-reflection consistently outperform those that do not take cross-reflection into account.
Third, similar to the findings in LLM-Rec, our method, which leverages augmented information from the cross-reflection prompting components, achieves superior performance compared to all other baselines.
MovieLens-1M Amazon-Software Precision@10 Recall@10 NDCG@10 Precision@10 Recall@10 NDCG@10 Item Popularity 0.0426 0.0428 0.0530 0.0082 0.0777 0.0571 Text 0.2877 0.2424 0.3597 0.0231 0.2175 0.1568 KAR (Xi et al. 2023) 0.2990 0.2575 0.3724 0.0189 0.1686 0.1175 LLM-Rec (Lyu et al. 2024) 0.2965 0.2547 0.3710 0.0424 0.3911 0.2843 Text + Image 0.2936 0.2529 0.3645 0.0268 0.2522 0.1966 Rec-GPT4V (Liu et al. 2024c) 0.3004 0.2589 0.3748 0.0254 0.2401 0.2000 CoT (Wei et al. 2022) 0.3039 0.2636 0.3806 0.0305 0.2859 0.2157 Visual-only prompting 0.3025 0.2626 0.3770 0.0274 0.2591 0.1988 Visual-textual prompting 0.3036 0.2633 0.3790 0.0267 0.2490 0.1914 Cross-reflection - separate 0.3060 0.2649 0.3830 0.0385 0.3537 0.2661 Cross-reflection - combined 0.3067 0.2681 0.3822 0.0318 0.2968 0.2414 X-Reflect 0.3111 0.2746 0.3900 0.0431 0.3969 0.3024 (+5.96) (+8.58) (+7.00) (+60.82) (+57.38) (+53.81)
Top-1 Recommendation Performance
MovieLens-1M Amazon-Software Precision@1 Recall@1 Precision@1 Recall@1 Text 0.4109 0.0401 0.1050 0.1017 KAR (Xi et al. 2023) 0.4248 0.0431 0.2108 0.2039 LLM-Rec (Lyu et al. 2024) 0.4150 0.0412 0.2191 0.2135 Text + Image 0.3966 0.0390 0.1550 0.1508 Rec-GPT4V (Liu et al. 2024c) 0.4115 0.0415 0.1863 0.1826 CoT (Wei et al. 2022) 0.4214 0.0430 0.1803 0.1744 Visual prompting 0.4103 0.0407 0.1892 0.1858 Visual-textual prompting 0.4131 0.0415 0.1813 0.1766 Cross-reflection - separate 0.4226 0.0424 0.2118 0.2034 Cross-reflection - combined 0.4119 0.0413 0.2080 0.2027 X-Reflect 0.4232 0.0436 0.2271 0.2177 (+6.71) (+11.79) (+46.52) (+44.36)
Table 3 displays the recommendation performance for Top-1 recommendations, measured by Precision@1 and Recall@1. These metrics assess the quality of recommendations when only one item is suggested. Since only one item is recommended, NDCG@1 simplifies to Precision@1. X-Reflect demonstrates strong performance in Top-1 recommendations, achieving the highest Precision@1 and Recall@1 on the Amazon-Software dataset, and the highest Recall@1 with the second-highest Precision@1 on MovieLens-1M. KAR (Xi et al. 2023) achieves the highest Precision@1 on MovieLens-1M, likely because it uses open-world knowledge to enrich item descriptions. For movies, unique characteristics that are better represented as knowledge may allow KAR to precisely augment special information, making it particularly effective in Top-1 recommendations. However, as seen in the Top-10 recommendation results, KAR’s performance diminishes, potentially due to its limited ability to enrich general information effectively.
Generalizability of the Proposed Framework
To assess the generalizability of our framework, we replace the GPT-4V backbone with two other widely used LMMs–InternLM and LLaVA–while keeping the experimental settings consistent across all baselines. Tables 5 and 4 show the corresponding recommendation performance. With LLaVA as the backbone, X-Reflect continues to outperform other prompting baselines. When InternLM is used as the backbone, our method remains one of the top-performing frameworks, achieving the highest Precision@10 across both datasets and excelling on the Amazon-Software dataset.
MovieLens-1M Amazon-Software Precision@10 Recall@10 NDCG@10 Precision@10 Recall@10 NDCG@10 Text + Image 0.2936 0.2529 0.3645 0.0268 0.2522 0.1966 Rec-GPT4V (Liu et al. 2024c) 0.2946 0.2523 0.3666 0.0309 0.2916 0.2245 CoT (Wei et al. 2022) 0.3001 0.2603 0.3746 0.0255 0.2406 0.2033 Visual-only prompting 0.2958 0.2542 0.3689 0.0313 0.2939 0.2388 Visual-textual prompting 0.2998 0.2570 0.3740 0.0317 0.2962 0.2380 Cross-reflection - separate 0.2909 0.2464 0.3635 0.0438 0.4048 0.3257 Cross-reflection - combined 0.3002 0.2573 0.3752 0.0278 0.2602 0.2210 X-Reflect 0.3002 0.2591 0.3736 0.0457 0.4172 0.3232 (+2.25) (+2.45) (+2.50) (+70.52) (+65.42) (+64.39)
MovieLens-1M Amazon-Software Precision@10 Recall@10 NDCG@10 Precision@10 Recall@10 NDCG@10 Text + Image 0.2936 0.2529 0.3645 0.0268 0.2522 0.1966 Rec-GPT4V (Liu et al. 2024c) 0.3002 0.2569 0.3756 0.0320 0.3039 0.2481 CoT (Wei et al. 2022) 0.3007 0.2574 0.3760 0.0288 0.2702 0.2106 Visual-only prompting 0.2968 0.2544 0.3700 0.0292 0.2747 0.2184 Visual-textual prompting 0.3005 0.2578 0.3754 0.0309 0.2884 0.2341 Cross-reflection - separate 0.3017 0.2609 0.3765 0.0398 0.3696 0.2970 Cross-reflection - combined 0.2976 0.2558 0.3715 0.0317 0.2969 0.2473 X-Reflect 0.3026 0.2640 0.3783 0.0411 0.3829 0.3049 (+3.07) (+4.39) (+3.79) (+53.36) (+51.82) (+55.09)
Further Discussion on Augmented Content
To better understand the impact of cross-reflection, we compute the cosine similarity between the embeddings of the original text description and the image description, as well as the responses generated through text-only prompting (LLM-Rec), multimodal prompting (CoT), and cross-reflection (separate) multimodal prompting.
For each item in the MovieLens-1M dataset, we convert its text description and image description into embeddings, defining the cosine similarity between these two embeddings as reference similarity. A higher value indicates greater semantic similarity between the text and image descriptions, serving as a proxy for how much additional information the image description contains compared to the original text description.
Similarly, we compute the cosine similarity between the text description embeddings and the responses generated through text-only prompting, multimodal prompting, and cross-reflection multimodal prompting. We then compare these cosine similarities with the reference similarity. The results are presented in Figure 3.
In each subplot, the x-axis represents the cosine similarity between the embeddings of the text and image descriptions, while the y-axis shows the cosine similarity between the embeddings of the text description and the responses generated through the other three prompting baselines.
The key finding is that cross-reflection multimodal prompting generates semantically more diverse information compared to other prompting strategies and image descriptions. This aligns with our earlier discussion, where we note that visual information from item images contains additional details not present in the original text description. However, directly prompting LMMs to describe an image may not elicit a comprehensive understanding of the item.
Even combining text and image in the prompt is not sufficient; explicit instruction is necessary to direct LMMs to focus on non-overlapping information across modalities. Although not identical, a similar pattern has been observed in multi-input image prompting (Zhang et al. 2024), where explicit instructions to focus on similarities and differences among multiple image inputs improve multimodal reasoning in LMMs.
It is important to note that the utility of augmented information in downstream recommendation tasks is still under investigation. While this experiment demonstrates that semanticallt more diverse information often leads to better performance, this is not an absolute rule. As seen in the experiments, generating less granular responses can sometimes also enhance performance in downstream tasks. This remains an ongoing area of research in the recommendation domain, and we plan to explore this further in future work.
Ablation on Prompts
To assess the robustness of our prompting framework, we modify our prompts to instruct LMMs to generate only unique keywords while maintaining the explicit instruction to identify whether the text and image information is supportive or contradictory. The results are presented in Table 6. This variant, termed X-Reflect-keyword, shows that X-Reflect is robust across different prompt designs. The performance remains consistent with the inclusion of explicit cross-reflection instructions.
MovieLens-1M Amazon-Software Precision@10 Recall@10 NDCG@10 Precision@10 Recall@10 NDCG@10 X-Reflect 0.3111 0.2746 0.3900 0.0431 0.3969 0.3024 X-Reflect-keyword 0.3128 0.2754 0.3908 0.0425 0.3917 0.3125
Discussions and Conclusions
In this paper, we introduce a novel framework, X-Reflect, which leverages cross-reflection prompting to enrich item descriptions by integrating both textual and visual modalities. Through a series of comprehensive experiments on two widely recognized recommendation benchmarks–MovieLens-1M and Amazon-Software–we demonstrate that our approach consistently outperforms several baseline prompting strategies, particularly those that do not incorporate cross-reflection.
Our findings suggest that by prompting Large Multimodal Models (LMMs) to explicitly identify and reconcile supportive or conflicting information between text and image, X-Reflect is able to generate more diverse and contextually rich item representations. This, in turn, enhances the effectiveness of recommendation systems by providing a more holistic understanding of the items being recommended.
Limitations
However, the proposed framework is not without its limitations. One notable drawback is the increased computational overhead associated with the need for multiple LMM inferences per item. This requirement can be resource-intensive, especially when processing large-scale datasets or deploying the system in real-time applications. To mitigate this issue, future research could explore optimization strategies, such as selectively applying cross-reflection prompting to a subset of items that are likely to benefit the most from this process. Additionally, advances in model efficiency, such as improved inference algorithms or more lightweight LMM architectures, could further reduce the computational burden.
Despite these challenges, the potential of cross-reflection prompting to enhance multimodal recommendations is significant. By continuing to refine and optimize this approach, we believe it can be a valuable tool in the ongoing effort to develop more accurate, personalized, and context-aware recommendation systems. Future work will also explore the integration of supervised fine-tuning and continuous pre-training to further improve the quality of responses generated by LMMs, as well as investigating the types of augmented information that are most beneficial for various downstream tasks.
Ethical Considerations
Recommendation systems, particularly those utilizing multimodal prompting methods like ours, significantly influence user behavior. There is a risk that such systems could inadvertently promote biased or harmful content or products, especially if the context of both visual and textual data is not properly assessed. For instance, our method, which integrates information from both images and text, might recommend items that are visually appealing but are contextually inappropriate or unsafe if the nuances of the data are misinterpreted.
To mitigate these risks, it is crucial to incorporate robust content filtering mechanisms within the multimodal framework to detect and block the promotion of harmful or inappropriate content. This includes not only evaluating the textual data for potential issues but also ensuring that the visual data aligns with safety standards. Additionally, our system should be designed with a strong emphasis on user well-being, avoiding recommendations that could lead to negative outcomes, such as encouraging excessive spending, unhealthy behaviors, or exposure to harmful content. By prioritizing safety and ethical considerations in our multimodal prompting approach, we can better protect users and promote positive interactions with the recommendation system.
Appendix A Appendix
KAR Augmentation Details
In the work by Xi et al. (2023), a specific prompt was designed to elicit factual knowledge about items. The prompt instructed the model to: “Introduce {item}, {item description} and describe its attributes precisely (including but not limited to {scenario-specific factors})”. The {scenario-specific factors} were identified by prompting an LLM with the following instruction: “List the importance factors or features that determine whether a user will be interested in a {item}.”
For the MovieLens-1M dataset, we directly use the factual knowledge about movies generated in their study. For the Amazon-Software dataset, we apply the same approach to first identify the {scenario-specific factors}. The prompt was adapted to: “List the importance factors or features that determine whether a user will be interested in software.” The set of factors generated by ChatGPT includes functionality, user interface, user experience, performance, compatibility, cost, security, support and documentation, customizability, scalability, reliability, updates and maintenance, reputation and reviews, and accessibility.
These identified factors are then used in prompts to enrich the factual knowledge of the items in the Amazon-Software dataset.
Example Responses
The example full response generated through visual-only prompting in Figure 1 is as follows:
“This is a movie poster for "One Flew Over the Cuckoo’s Nest." The poster features a black and white image of a man looking upwards with a smile on his face. He is wearing a beanie hat and a leather jacket. Above him, the title of the movie is written in bold, capitalized letters, with a graphic of a goose flying over the lettering. The name "JACK NICHOLSON" is prominently displayed at the top of the poster.
Below the image of the man, there is additional text providing information about the film, including the production company "Fantasy Films," the fact that it is a Milos Forman film, and other cast and crew details. The bottom of the poster also indicates that the movie is based on the novel by Ken Kesey, and it mentions the availability of the novel in paperback. The United Artists logo is visible at the bottom right corner, along with a "Restricted" audience rating symbol on the left side. ”
The example full response generated through visual-textual prompting in Figure 1 is as follows:
“The item you’re referring to is a promotional poster for the film "One Flew Over the Cuckoo’s Nest," starring Jack Nicholson. The poster features an image of Jack Nicholson looking up, with a smile on his face, wearing a beanie hat. The title of the film is prominently displayed above his image, with a graphic of a goose flying over the text, which plays on the title’s metaphor. The poster also includes the names of other key contributors to the film, such as Louise Fletcher and William Redfield, as well as the director, Milos Forman, and producers Saul Zaentz and Michael Douglas. The film is based on the novel by Ken Kesey, and the screenplay was written by Lawrence Hauben and Bo Goldman.
The description you provided outlines the central plot of the movie, where the main character, portrayed by Jack Nicholson, pretends to be insane to avoid prison labor, but his plan backfires when he is sent to a mental institution. There, he encounters a tyrannical head nurse and faces new challenges. The film is known for its critical look at the institutional processes and the human spirit’s struggle against oppression. It is a classic film that has been celebrated for its performances, direction, and powerful storytelling. ”
The example full response generated through cross-reflection prompting in Figure 1 is as follows:
“The item description and the image description provided are mostly supportive of each other, but there are some notable differences.
Supportive Elements:
Character Description: The item description mentions a "criminal" who ends up in a mental institution. The image shows a man (presumably the protagonist, played by Jack Nicholson) with a relaxed, slightly mischievous smile, which could suggest someone who is not entirely compliant or is defying the situation, aligning with the storyline of a character challenging the system.
Title and Actor Name: The image description and item description both correctly reference the title of the movie, "One Flew Over the Cuckoo’s Nest," and mention Jack Nicholson, who is the lead actor in the film.
Discrepancies:
Image Context vs. Plot: The image description focuses on the man smiling and looking upwards, which might suggest a lighter or more hopeful moment. However, the item description suggests a darker and more intense story about a criminal struggling against an oppressive system in a mental institution. The tone of the image (relaxed, slightly cheerful) seems to conflict with the intense and serious tone suggested by the item description.
Graphic Representation: The image description mentions a "graphic of a goose flying over the lettering," which could symbolize the idea of escape or freedom, tying into the plot about a character seeking freedom from oppression. However, this symbolic element is not discussed in the item description.
Conclusion:
While the image and item description are generally aligned, there is a slight tonal discrepancy where the image suggests a more upbeat or defiant mood, whereas the item description suggests a more serious, oppressive environment. This could lead to different interpretations of the movie’s theme if solely relying on one of the descriptions. The differences are not major but might slightly alter a viewer’s initial expectations.”